OpenStudy (anonymous):

Please help! A landing airplane accelerates at -3.10 m/sec^2 down a 3.80 km runway. What is the fastest speed that the pilot could land and still stop at the end of the runway?

OpenStudy (anonymous):

\[x = vt + \frac{1}{2}at^{2}\]

OpenStudy (anonymous):

sorry that doesnt help give me a second

OpenStudy (anonymous):

I am only given the acceleration and the distance... and I think I need to find the initial velocity

OpenStudy (anonymous):

∆d = Vo*t + 1/2at^2.... maybe use this equation?

OpenStudy (anonymous):

yeah exactly, you want an equation without time

OpenStudy (anonymous):

Ahhh... V^2 = Vo^2 + 2a∆d

OpenStudy (anonymous):

\[v^{2}=u^{2} + 2ax\]?

OpenStudy (anonymous):

yeah that one